OpenStudy (anonymous):

Help Me Please !! Simplify the expression. 3(x + 3) – 2x + 4 A. 5x + 5 B. x + 13 C. 6x – 2 D. 3x – 11

OpenStudy (sid123):

3(x+3)-2x+4=3X+9-2x+4 =3x-2x+9+4 =x+13

OpenStudy (anonymous):

Thank You !